# Detecting Misbehaviors of Large Vision-Language Models by Evidential Uncertainty Quantification

### v1

@@ 5 Detection Performance Analysis
Hallucination datas from (Li et al., 2023) and (Wu et al., 2024a).
@@
Jailbreak attacks from (Gong et al., 2025), (Li et al., 2024), (Qi et al., 2024), and (Goh et al., 2021).
@@ 5.2 Ablation Study
Prompting for Abstention We tested external prompting by adding a "None of the above" option (Wang and Nalisnick, 2025). As shown in Table 5, the models rarely selected it. Qwen chose it only 0.27% of the time, and Intern 0.00%. Even after reinforcing the instruction with If you are unsure, please select "None of the above"., the selection rates remained extremely low, with Qwen choosing it 4.39% of the time and Intern 0.53%, indicating persistent overconfidence in multiple-choice scenarios. This shows that prompt-based strategies have limited ability to elicit uncertainty, underscoring the need for methods that do not rely on prompting.
@@ 6 Discussion
Future work
A promising direction is to extend our framework to black-box settings by estimating uncertainty from final outputs would enable application to APIs. Additionally, uncertainty signals could be integrated into the generation process itself, allowing models to self-correct during reasoning.
@@ 7 Conclusion
8 Ethics Statement
@@ A.7.1 Datasets
In hallucination scenarios, evaluation is conducted on POPE (Li et al., 2023) and R-Bench (Wu et al., 2024a), respectively, targeting object and relation hallucinations.
We follow the evaluation protocol proposed in POPE (Li et al., 2023), which formulates object hallucination detection as a binary (Yes-or-No) task. Built on the MS COCO validation set (Lin et al., 2014), POPE prompts LVLMs with queries such as Is there a chair in the image? and expects definitive binary responses. This task design enables precise assessment of hallucination by systematically querying for nonexistent objects. To increase the likelihood of hallucination, we adopt POPE’s Adversarial Sampling strategy: for each image, objects are ranked by their co-occurrence frequency with ground-truth objects, and the top-$k$ frequently co-occurring but absent objects are selected as hallucination probes. We further evaluate relational hallucination using the R-Bench benchmark (Wu et al., 2024a), which is specifically designed to assess the ability of LVLMs to correctly perceive and reason about object relationships. R-Bench includes two types of binary (Yes-or-No) questions: image-level and instance-level. The image-level questions probe whether a particular relationship exists anywhere in the image, while the instance-level questions target relationships between specific object instances, denoted by color-coded bounding boxes or masks. The latter is particularly useful for assessing fine-grained and localized relational understanding. All examples are constructed from the validation split of the nocaps dataset (Agrawal et al., 2019), and we follow the data preparation procedure provided in the official repository. [^4] https://github.com/mrwu-mac/R-Bench
@@
For jailbreak scenarios, we evaluate a range of jailbreak attacks, including FigStep (Gong et al., 2025), Hades (Li et al., 2024), and VisualAdv (Qi et al., 2024), all implemented via the Jailbreak Benchmark [^5] https://github.com/thunxxx/MLLM-Jailbreak-evaluation-MMJ-bench. We further simulate typographic attacks following the protocol of (Goh et al., 2021).
@@
  • For Hades (Li et al., 2024) and VisualAdv (Qi et al., 2024), we utilize the harmful instructions as provided by the Jailbreak Benchmark.
@@ A.8.3 Single-Modality Evaluation of Evidential Uncertainty
To further demonstrate the adaptability of our method, we conducted additional experiments on single-modality models. We performed a controlled experiment using a LeNet classifier trained on the handwritten digits dataset MNIST (LeCun, 1998) and the German Traffic Sign Recognition Benchmark (GTSRB) (Stallkamp et al., 2011), with FashionMNIST (Xiao et al., 2017) serving as out-of-distribution (OOD) data and FGSM-generated adversarial examples (Goodfellow et al., 2014). For comparison, we employed several classical uncertainty quantification methods: MC Dropout (Gal and Ghahramani, 2016) (100 iterations), Deep Ensembles (Lakshminarayanan et al., 2017) (5 models), and Evidential Deep Learning (EDL) (Sensoy et al., 2018). As shown in Table 11, $\mathrm{CF}$ and $\mathrm{IG}$ achieve competitive or superior performance compared to these baselines, attaining high AUROC scores for both adversarial and OOD detection tasks.

### v2

@@ 5 Detection Performance Analysis
Hallucination datas from (Li et al., 2023b) and (Wu et al., 2024a).
@@
Jailbreak attacks from (Gong et al., 2025), (Li et al., 2024b), (Qi et al., 2024), and (Goh et al., 2021).
@@ 5.2 Ablation Study
Prompting for Abstention We tested external prompting by adding a "None of the above" option (Wang & Nalisnick, 2025). As shown in Table 5, the models rarely selected it. Qwen chose it only 0.27% of the time, and Intern 0.00%. Even after reinforcing the instruction with If you are unsure, please select "None of the above"., the selection rates remained extremely low, with Qwen choosing it 4.39% of the time and Intern 0.53%, indicating persistent overconfidence in multiple-choice scenarios. This shows that prompt-based strategies have limited ability to elicit uncertainty, underscoring the need for methods that do not rely on prompting.
@@ 7 Conclusion
Acknowledgments
This work was supported by the National Key Research and Development Program (Grant No. 2024YFE0202900), the National Natural Science Foundation of China (Grant Nos. 62436001, 62536001, and 62406021), and the Joint Foundation of Ministry of Education for Innovation team (Grant No. 8091B042235).
Ethics Statement
@@ A.7.1 Datasets
In hallucination scenarios, evaluation is conducted on POPE (Li et al., 2023b) and R-Bench (Wu et al., 2024a), respectively, targeting object and relation hallucinations.
We follow the evaluation protocol proposed in POPE (Li et al., 2023b), which formulates object hallucination detection as a binary (Yes-or-No) task. Built on the MS COCO validation set (Lin et al., 2014), POPE prompts LVLMs with queries such as Is there a chair in the image? and expects definitive binary responses. This task design enables precise assessment of hallucination by systematically querying for nonexistent objects. To increase the likelihood of hallucination, we adopt POPE’s Adversarial Sampling strategy: for each image, objects are ranked by their co-occurrence frequency with ground-truth objects, and the top-$k$ frequently co-occurring but absent objects are selected as hallucination probes. We further evaluate relational hallucination using the R-Bench benchmark (Wu et al., 2024a), which is specifically designed to assess the ability of LVLMs to correctly perceive and reason about object relationships. R-Bench includes two types of binary (Yes-or-No) questions: image-level and instance-level. The image-level questions probe whether a particular relationship exists anywhere in the image, while the instance-level questions target relationships between specific object instances, denoted by color-coded bounding boxes or masks. The latter is particularly useful for assessing fine-grained and localized relational understanding. All examples are constructed from the validation split of the nocaps dataset (Agrawal et al., 2019), and we follow the data preparation procedure provided in the official repository. [^4] https://github.com/mrwu-mac/R-Bench
@@
For jailbreak scenarios, we evaluate a range of jailbreak attacks, including FigStep (Gong et al., 2025), Hades (Li et al., 2024b), and VisualAdv (Qi et al., 2024), all implemented via the Jailbreak Benchmark [^5] https://github.com/thunxxx/MLLM-Jailbreak-evaluation-MMJ-bench. We further simulate typographic attacks following the protocol of (Goh et al., 2021).
@@
  • For Hades (Li et al., 2024b) and VisualAdv (Qi et al., 2024), we utilize the harmful instructions as provided by the Jailbreak Benchmark.
@@ A.8.3 Single-Modality Evaluation of Evidential Uncertainty
To further demonstrate the adaptability of our method, we conducted additional experiments on single-modality models. We performed a controlled experiment using a LeNet classifier trained on the handwritten digits dataset MNIST (LeCun, 1998) and the German Traffic Sign Recognition Benchmark (GTSRB) (Stallkamp et al., 2011), with FashionMNIST (Xiao et al., 2017) serving as out-of-distribution (OOD) data and FGSM-generated adversarial examples (Goodfellow et al., 2014). For comparison, we employed several classical uncertainty quantification methods: MC Dropout (Gal & Ghahramani, 2016) (100 iterations), Deep Ensembles (Lakshminarayanan et al., 2017) (5 models), and Evidential Deep Learning (EDL) (Sensoy et al., 2018). As shown in Table 11, $\mathrm{CF}$ and $\mathrm{IG}$ achieve competitive or superior performance compared to these baselines, attaining high AUROC scores for both adversarial and OOD detection tasks.